OpenStudy (anonymous):

Find the inverse of the following matrix. 7 3 4 2

OpenStudy (anonymous):

\[\frac{1}{(7*2)-(3*4)}\]\[\frac{1}{14-12}=\frac{1}{2}\]

OpenStudy (anonymous):

Do you know How to find Co-Factors ?

OpenStudy (anonymous):

Do u want the ANSWER ? or Should I walk u through the Problem..

OpenStudy (anonymous):

Please walk me through the problem. I have no idea where to begin.

OpenStudy (anonymous):

Ok

OpenStudy (salina):

\[\frac{1}{(a*d)-(b*c)}\]time [d -b] [-c a]

OpenStudy (salina):

\[\frac{1}{(7*2)-(4*3)}\]=1/2

OpenStudy (anonymous):

a b c d to find the inverse find the determinant (ad-bc) The inverse is 1/(ad-bc) times the matrix d -b -c a so in this case it is 1/2 [ 2 -3 -4 7 ] =[ 1 -1.5 -2 3.5 ]

OpenStudy (salina):

1/2 * [2 -3] [1 -3/2] [-4 7] = [-2 7/2]

OpenStudy (anonymous):

Thank you very much!

OpenStudy (anonymous):

very easy way is \[\left[\begin{matrix}A & B \\ C & D\end{matrix}\right]\] take the determinate 1/(AD-CB) and flip C and B and negate A and D

OpenStudy (anonymous):

swap A and D and negate B and C